# Bayesian Low-Rank LeArning (Bella): A Practical Approach to Bayesian Neural Networks

### v1

@@ 4 Bayesian Low-Rank Learning (Bella) for SVGD
$\mathbf{A}_{i}=\mathbf{A}_{i}-\epsilon_{i}\sum_{j=1}^{n}\hat{\boldsymbol{\phi}}{}_{j}^{*}(\mathbf{A}_{i}),\quad\mathbf{B}_{i}=\mathbf{B}_{i}-\epsilon_{i}\sum_{j=1}^{n}\hat{\boldsymbol{\phi}}{}_{j}^{*}(\mathbf{B}_{i})\text{\quad with}$ (4)
$\hat{\boldsymbol{\phi}}{}_{j}^{*}(\mathbf{B}_{i})$ $=k\left(\mathbf{B}_{j}\mathbf{A}_{j},\mathbf{B}_{i}\mathbf{A}_{i}\right)\!\nabla_{\mathbf{B}_{i}}\ell\left(f(\mathbf{x};\boldsymbol{\theta}_{0}+\mathbf{B}_{i}\mathbf{A}_{i}),y\right)-\frac{\gamma}{n}\nabla_{\mathbf{B}_{i}}k\left(\mathbf{B}_{j}\mathbf{A}_{j},\mathbf{B}_{i}\mathbf{A}_{i}\right),$
$\hat{\boldsymbol{\phi}}{}_{j}^{*}(\mathbf{A}_{i})$ $=k\left(\mathbf{B}_{j}\mathbf{A}_{j},\mathbf{B}_{i}\mathbf{A}_{i}\right)\!\nabla_{\mathbf{A}_{i}}\ell\left(f(\mathbf{x};\boldsymbol{\theta}_{0}+\mathbf{B}_{i}\mathbf{A}_{i}),y\right)-\frac{\gamma}{n}\nabla_{\mathbf{A}_{i}}k\left(\mathbf{B}_{j}\mathbf{A}_{j},\mathbf{B}_{i}\mathbf{A}_{i}\right)\,.$
Here, we have placed the (improper) uniform prior on $(\mathbf{A}_{i},\mathbf{B}_{i})$, but other choices are possible. Uniform prior implies that we only require the gradients of the log-likelihood computed on a mini-batch of data, as well as the kernel function to encourage particle diversity. Note that the kernel function on $(\mathbf{A}_{i},\mathbf{B}_{i})$ is given by $k\left(\boldsymbol{\theta}_{0}+\mathbf{B}_{j}\mathbf{A}_{j},\boldsymbol{\theta}_{0}+\mathbf{B}_{i}\mathbf{A}_{i}\right)$, which ensures that the similarity is computed on the original parameter space, and in the commonly used case of shift-invariant kernels, this simplifies to $k\left(\mathbf{B}_{j}\mathbf{A}_{j},\mathbf{B}_{i}\mathbf{A}_{i}\right)$. Further simplifications are obtained for specific kernel functions – in particular, in the case of Gaussian RBF, while the naive implementation would require the cost of $O(rd_{1}d_{2})$ for a single kernel evaluation, we can bring it down to $O(r^{2}(d_{1}+d_{2}))$ using standard trace manipulation, as described in the appendix. This procedure can be repeated across all dense layers.
@@ 5 Empirical Experiments and Results
In this section, we provide an in-depth overview of our experimental setup, detailing the methodology, equipment, and procedures employed in the implementation of our Bella. We meticulously outline the configuration settings, dataset, and the criteria used for evaluating outcomes, ensuring a transparent and reproducible framework. Following this, we present the experimental results, offering a thorough analysis of the data obtained, including statistical evaluations and interpretations of the findings in the context of our research objectives. This comprehensive approach ensures a clear understanding of the experiment’s execution and its contributions to the field.
@@ 5.1 Experimental Set-up
Datasets. In this research, we have employed a variety of datasets, each selected for its relevance and contribution to the field of image recognition and computer vision. These datasets include CIFAR-10, CIFAR-100 [37], CIFAR-10-C [38], STL-10 [39], CAMELYON17 [40], ImageNet [35], and DomainNet [41]. We also employ VQA v2 dataset containing billions of questions answers given hundred thousands of images [42], which is utilized in Visual Question Answering (VQA) task. Detailed information about datasets is presented in Appendix H and Appendix F.
Networks. In our experiments, we employed the CLIP ViT-B/32 model [43], a pre-trained variant utilizing contrastive supervision from image-text pairs, as initially introduced in the seminal CLIP research [44]. We conducted end-to-end fine-tuning, adjusting all model parameters, a strategy typically yielding higher accuracy compared to training only the final linear layer. In our methodology, we selectively fine-tune only the image encoder of the CLIP model while maintaining the text encoder frozen as it is common [45]. To elaborate, both during training and inference phases, the final output is derived through an inner product between the output produced by the image encoder and the embedding generated by the frozen text encoder. For both our ensemble-based methods which include both traditional ensemble [22] and the Stein Variational Gradient Descent (SVGD) approach [25], we employed the technique of utilizing the logits (unnormalized outputs) of the models, as described in [46]. This strategy has been consistently applied in our ensemble configurations, encompassing traditional ensembles as well as SVGD. We also employ Ensemble networks as it was shown to be a good and efficient way to approximate the posterior distribution to be baselines [21]. For VQA task, we employ the SoTA LLaVA-1.5 7B with billions of parameters to showcase the effectiveness of our Bella on large-scale network architecture. Details regarding hyper-parameters are in Appendix H.
5.2 Benchmark Task Performance and Cost Efficiency
The performance comparison of our Bella models with their respective base models, as well as with Vision Bayesian Lora (VBL)—a derived Laplacian sampling method from [20] for vision tasks—is delineated in Table 1. The respective costs comparison on training (in terms of trainable parameters) is shown in Table 2—notably, the grey columns are the ones used for generating the results in Table 1.
@@ 5.2 Benchmark Task Performance and Cost Efficiency
Ensemble (n=5) SVGD (n=5) Single VBL Ensemble (n=5) SVGD (n=5)
CIFAR10 97.57 97.95 97.20 94.41 97.54 97.87
CIFAR100 86.50 88.09 86.17 86.22 86.89 87.32
CAMELYON17 94.47 94.89 92.38 92.96 94.23 95.21
DomainNet 82.70 84.47 74.61 82.97 85.07 85.42
ImageNet 77.29 78.24 76.87 - 78.93 79.36
Table 1: Comparing performance of Bella models with their baseline (Base) counterparts on vision benchmarks. The results confirm Bella models are well-performing and on par with their baseline Ensemble, SVGD, and VBL whilst consuming only a fraction of cost (see Table 2 grey columns for corresponding costs). The best-performing results are in bold, the second-best in underline, and the least favorable are in italic for emphasis.
                         Models SVGD base SVGD Bella Single
@@
Table 2: Computational cost to train different models based on CLIP architecture for different datasets. Notably, with SVGD we can only train up to $n$=40 particles on a A6000 48 GB GPU, while we can increase to more than 100 parameter particles with our Bella with negligible increase of GPU consumption.
@@
Notably, in Table 1 with Bella SVGD, with 1.6% of the trainable parameters in comparison to the Single baseline, leads to approximately 2% and 10% increase in performance on the OOD tasks of DomainNet and CAMELYON17, respectively; whilst achieving comparable performance with the current SVGD implementation (SVGD baseline model).
@@
Along with Table 1, Table 2 demonstrates the primary benefit of our approach—the significant reduction in memory and storage needs. For example, Table 2, we obtain approximately a 5$\times$ reduction in model size—that is from 2,222 MB for a 5 particle SVGD baseline model to just 440 MB for a Bella SVGD model with 5 particles. This efficiency not only reduces the demands on GPUs but also minimizes potential I/O bottlenecks.
@@
Moreover, the reduced GPU demand, as shown in Table 2, facilitates larger mini-batch sizes during training to speedup the training process. More crucially, it enables the enhancement of the Bayesian posterior’s parameter particles to over 100, a feat previously unattainable with current SVGD implementations. Significantly, constructing a 100 parameter Bayesian approximation consumes only 5.19 GB memory compared to current SVGD implementations for Bayesian models (SVGD base) needing over 6 GB for even a 3 particle approximation.
@@
Interestingly, our on par results of Bella models with current SVGD and ensemble approximations of the posterior provide empirical evidence that with constrained model parameters, it is still possible to reach the diverse modes of the posterior. Further, the results support recent conjectures on mode connectivity [10].
Models DomainNet Dataset Average
@@
Real Clip-Art Infograph Paint Sketch
Single base 74.61 55.29 31.81 53.87 43.84 51.89
VBL 82.97 59.94 26.56 52.81 46.96 53.85
Ensemble Bella 82.70 61.22 28.15 55.32 51.58 55.60
Ensemble base 85.07 65.40 36.07 57.90 54.22 59.73
SVGD Bella 84.47 63.67 32.22 56.83 54.86 58.41
SVGD base 85.42 65.53 36.58 58.18 55.47 60.24
Table 3: Out-of-distribution performance of our method Bella on the DomainNet benchmark. Bella models perform comparable to current SVGD and Ensemble, whilst performing significantly better than the Single base. (Recall Bayesian approximation are from $n=5$ particles)
5.3 Performance on Out-Of-Disrtibution (OOD) Datasets
@@
Assessing the robustness of machine learning systems to unseen conditions is crucial, especially their ability to generalize to out-of-distribution (OOD) data. We evaluate robutness to ODD using multiple OOD benchmarks.
@@
First, we use the DomainNet dataset, one of the most diverse domain adaptation datasets, covering a wide range of visual styles from real images to abstract art. This variety provides a challenging test bed for algorithms aiming to bridge different visual domains. Our study involves training the CLIP network on DomainNet’s ‘Real’ subset and testing its generalization across various domains (see Table 3). Second, CIFAR-10-C, a corrupted version of CIFAR-10, helps assess model generalization and robustness (see Table 4). Third, the STL-10 dataset, with significant label overlap with CIFAR-10, serves as a relevant OOD test case (see Appendix B).
@@
Bella models, with significantly better efficiency, demonstrate competitive performance against more resource-intensive implementations with Ensemble and SVGD baselines across both DomainNet and CIFA-10-C benchmarks. All of the Bayesian approximations, including our scalable and efficient method, outperform the Single model baseline.

### v2

@@ 4 Bayesian Low-Rank Learning (Bella)
$\mathbf{A}_{i}=\mathbf{A}_{i}-\epsilon_{i}\sum_{j=1}^{n}\hat{\boldsymbol{\phi}}{}_{j}^{*}(\mathbf{A}_{i}),~{}~{}~{}\mathbf{B}_{i}=\mathbf{B}_{i}-\epsilon_{i}\sum_{j=1}^{n}\hat{\boldsymbol{\phi}}{}_{j}^{*}(\mathbf{B}_{i})$
$\text{with\quad}\hat{\boldsymbol{\phi}}{}_{j}^{*}(\mathbf{B}_{i})$ $=k_{i,j}\!\nabla_{\mathbf{B}_{i}}p\left(y\mid\mathbf{x},\boldsymbol{\theta}_{0}+\mathbf{B}_{i}\mathbf{A}_{i}\right)-\frac{\gamma}{n}\nabla_{\mathbf{B}_{i}}k_{i,j},$
$\hat{\boldsymbol{\phi}}{}_{j}^{*}(\mathbf{A}_{i})$ $=k_{i,j}\!\nabla_{\mathbf{A}_{i}}p\left(y\mid\mathbf{x},\boldsymbol{\theta}_{0}+\mathbf{B}_{i}\mathbf{A}_{i}\right)-\frac{\gamma}{n}\nabla_{\mathbf{A}_{i}}k_{i,j}\,,$
where we denote $k_{i,j}=k\left(\mathbf{B}_{j}\mathbf{A}_{j},\mathbf{B}_{i}\mathbf{A}_{i}\right)$. Here, we have placed a zero-mean Gaussian prior on $(\mathbf{A}_{i},\mathbf{B}_{i})$, but other choices are possible. Note that the kernel function on $(\mathbf{A}_{i},\mathbf{B}_{i})$ is given by $k\left(\boldsymbol{\theta}_{0}+\mathbf{B}_{j}\mathbf{A}_{j},\boldsymbol{\theta}_{0}+\mathbf{B}_{i}\mathbf{A}_{i}\right)$, which ensures that the similarity is computed on the original parameter space, and in the commonly used case of shift-invariant kernels, this simplifies to $k\left(\mathbf{B}_{j}\mathbf{A}_{j},\mathbf{B}_{i}\mathbf{A}_{i}\right)$. Further simplifications are obtained for specific kernel functions – in particular, in the case of Gaussian RBF, while the naive implementation would require the cost of $O(rd_{1}d_{2})$ for a single kernel evaluation, we can bring it down to $O(r^{2}(d_{1}+d_{2}))$ using standard trace manipulation, as described in the Appendix G. This procedure can be repeated across all dense layers.
@@ 5 Empirical Experiments and Results
In this section, we provide an in-depth overview of our experimental setup, detailing the methodology, equipment, and procedures employed in the implementation of Bella. We aim to compare Bella with established ensemble-based methods and SVGD, both of which have demonstrated effectiveness in BDLs (Seligmann et al. 2024; Lakshminarayanan, Pritzel, and Blundell 2017).
@@ 5.1 Experimental Set-up
Datasets. In this research, we have employed a variety of datasets, each selected for their relevance and contribution, they include CIFAR-10, CIFAR-100 (Krizhevsky, Hinton et al. 2009), CIFAR-10-C (Hendrycks and Dietterich 2019), STL-10 (Coates, Ng, and Lee 2011), CAMELYON17 (Bandi et al. 2018), ImageNet (Russakovsky et al. 2015), and DomainNet (Peng et al. 2019). We also consider VQA v2 dataset utilized for Visual Question Answering (VQA). Additional details are in Appendix I and Appendix H.
@@
Ensemble (n=5) SVGD (n=5) Ensemble (n=5) SVGD (n=5) Single VBL
CIFAR10 97.32 $\pm$ 0.37% 97.57 $\pm$ 0.38% 97.26 $\pm$ 0.28% 97.56 $\pm$ 0.31% 96.86 $\pm$ 0.43% 94.04 $\pm$ 0.37%
CIFAR100 86.02 $\pm$ 0.48% 87.63 $\pm$ 0.46% 86.65 $\pm$ 0.24% 87.03 $\pm$ 0.29% 85.14 $\pm$ 1.3% 85.01 $\pm$ 1.21%
CAMELYON17 93.11 $\pm$ 1.36% 93.61 $\pm$ 1.28% 93.29 $\pm$ 0.94% 93.98 $\pm$ 1.23% 90.25 $\pm$ 2.31% 91.58 $\pm$ 1.38%
DomainNet 80.34 $\pm$ 2.36% 81.41 $\pm$ 3.06% 82.96 $\pm$ 2.11% 83.66 $\pm$ 1.76% 69.75 $\pm$ 4.86% 80.60 $\pm$ 2.37%
ImageNet 77.29 78.24 78.93 79.36 76.87 -
Figure 2: Left: Comparing performance of Bella models with their baseline (Base) counterparts on vision benchmarks. The results confirm Bella models are well-performing and on par with their baseline Ensemble, SVGD, and VBL whilst consuming only a fraction of cost (see Table 1 grey columns for corresponding costs). The best-performing results are in bold, the second-best in underline, and the least favorable are in italic for emphasis. Right: Bella achieves similar Accuracy with full SVGD (Base) only with a fraction of cost.
Neural Architecture. In our experiments, we employed the CLIP ViT-B/32 model (Ilharco et al. 2021), a pre-trained variant utilizing contrastive supervision from image-text pairs, as initially introduced in the seminal CLIP research (Radford et al. 2021). We conducted end-to-end fine-tuning of the image encoder, adjusting all model parameters, a strategy typically yielding higher accuracy than training only the final linear layer. For both our ensemble-based methods, we consider conventional ensemble and SVGD. We use the average logits (unnormalized outputs) to produce the output (Gontijo-Lopes, Dauphin, and Cubuk 2021). For VQA task, we employ the SoTA LLaVA-1.5-7B (Liu et al. 2023b) to showcase the effectiveness of Bella on large-scale network architecture. Details of the hyper-parameters are discussed in Appendix I.
5.2 Cost Efficiency
The accuracy comparison of our Bella models with their respective base models, as well as with Vision Bayesian Lora (VBL)—a derived Laplace’s approximation from (Yang et al. 2024) for vision tasks—is delineated in the left table in Figure 2. The respective costs comparison on training (in terms of trainable parameters) is shown in Table 1—notably, the grey columns are the ones used for generating the results in the left table in Figure 2.
                         Models Bseline Models (SVGD) Bella Models (SVGD) Single
@@ 5.2 Cost Efficiency
Table 1: Computational cost to train different models based on CLIP architecture for different datasets. Notably, with SVGD Baseline Models, we can only train up to $n$=40 particles on a A6000 48 GB GPU, while we can increase to more than 100 parameter particles with our Bella method with negligible increase of GPU consumption. The grey columns correspond to costs (parameters) of models in Figure 2.
@@
Notably, in Figure 2 with Bella SVGD, with 1.6% of the trainable parameters in comparison to the Single baseline, leads to approximately 2% and 10% increase in performance on the OOD tasks of DomainNet and CAMELYON17, respectively; whilst achieving comparable performance with the current SVGD implementation (SVGD baseline model).
@@
Along with Figure 2, Table 1 demonstrates the primary benefit of our approach—the significant reduction in memory and storage needs. For example, Table 1, we obtain approximately a 5$\times$ reduction in model size—that is from 2,222 MB for a 5 particle SVGD baseline model to just 440 MB for a Bella SVGD model with 5 particles. This efficiency not only reduces the demands on GPUs but also minimizes potential I/O bottlenecks.
@@
Moreover, the reduced GPU demand, as shown in Table 1, facilitates larger mini-batch sizes during training to speedup the training process. More crucially, it enables the enhancement of the Bayesian posterior’s parameter particles to over 100, a feat previously unattainable with current SVGD implementations. Significantly, constructing a 100 parameter Bayesian approximation consumes only 5.19 GB memory compared to current SVGD implementations for Bayesian models (SVGD base) needing over 6 GB for even a merely 3 particle approximation.
@@
Interestingly, our on-par results of Bella models with current SVGD and ensemble approximations of the posterior provide empirical evidence that with constrained model parameters, it is still possible to reach the diverse modes of the posterior. Further, the results support recent conjectures on mode connectivity (Gueta et al. 2023).
Models DomainNet Avg. CIFAR-10-C Avg.
@@
Real Clip-Art Infograph Paint Sketch Gaussian Blur Pixelate Spatter
Single base 74.61 55.29 31.81 53.87 43.84 51.89 90.58 77.94 92.14 78.53
VBL 82.97 59.94 26.56 52.81 46.96 53.85 89.34 76.43 89.21 70.76
Ensemble Bella 82.70 61.22 28.15 55.32 51.58 55.60 93.30 85.86 94.83 82.24
Ensemble base 85.07 65.40 36.07 57.90 54.22 59.73 91.92 86.49 93.45 84.61
SVGD Bella 84.47 63.67 32.22 56.83 54.86 58.41 94.05 88.70 94.85 83.77
SVGD base 85.42 65.53 36.58 58.18 55.47 60.24 93.02 86.75 95.19 86.84
Table 2: The out-of-distribution generalization performance of Bella, measured by accuracy (for number of particles $n=5$). Each column represents a specific shift, either real (DomainNet) or artificial (CIFAR-10-C). Further results can be found in Appendix D.
5.3 Out-Of-Disrtibution (OOD) Datasets
@@
Assessing the robustness of machine learning systems to unseen conditions is crucial, particularly regarding their ability to generalize to out-of-distribution (OOD) data. We evaluate robustness to OOD by utilizing multiple OOD benchmarks and estimating accuracy under various noise levels (distribution shifts) to further assess the performance of different baselines in these challenging scenarios.
@@
First, we use the DomainNet dataset, which is one of the most diverse domain adaptation datasets and spans a wide range of visual styles, from real images to abstract art. This variety provides a challenging test bed for algorithms aiming to bridge different visual domains. Our study involves training the CLIP network on the ‘Real’ subset of DomainNet and evaluating its generalization across various domains. Additionally, we use CIFAR-10-C, a corrupted version of CIFAR-10, to further assess the generalization and robustness of models trained on CIFAR-10 datasets. The results for both datasets are presented in LABEL:tab:domainnet_cifar10-c.
Furthermore, the STL-10 dataset, which has significant label overlap with CIFAR-10, serves as a relevant OOD test case for CIFAR-10 models (see Appendix B).
@@
Results from LABEL:tab:domainnet_cifar10-c and Appendix B show that Bella models, demonstrating significantly better efficiency, achieve competitive performance compared to more resource-intensive implementations with Ensemble and SVGD baselines across both DomainNet and CIFAR-10-C benchmarks. All Bayesian approximations, including our scalable and efficient method, outperform the single model baseline.
5.4 Comparing Uncertainty Estimations
Bayesian models capable of providing a theoretical basis for measuring model uncertainty. Also known as epistemic uncertainty, refers to uncertainty stemming from limitations in our knowledge or understanding of the underlying data generating process or the model itself. One of the ways to quantify model uncertainty is through mutual information estimates, following  (Gal et al. 2016).
Figure 3: Evaluation of uncertainty estimations using Mutual Information on CAMELYON17 and CIFAR-10-C datasets. $\uparrow$ MI for Misclassified Examples is better—denoted by the distribution shifting $\rightarrow$. In contrast, $\downarrow$ MI for Correctly Classified Examples is better—denoted by the distribution shifting $\leftarrow$. Notably, Single base models are not able to capture epistemic uncertainty (Gal et al. 2016).
5.5 Robustness against Adversarial Examples
In this section, we examine the resilience of our proposed Bella against adversarial attacks, specifically employing the $L_{\infty}$ Fast Gradient Sign Method (FGSM) across various attack budgets as detailed in Figure 4. This analysis aims to benchmark the robustness of our method in comparison to traditional models under adversarial conditions. We employ the robustness benchmark (Papernot et al. 2018) to deploy the attack on CIFAR-10 test set and report results in Figure 4.
Figure 4: Comparison of model robustness to $L_{\infty}$ FGSM adversarial attacks across varied attack budgets. The evaluation is on the CIFAR-10 dataset. Bella achieves comparable performance to current, baseline SVGD and Ensemble (number of particles is equal to $5$) implementations whilst demonstrating significantly better robustness (higher accuracy) than a Single baseline model.
5.6 Ablation Studies
This section undertakes a series of ablation studies to examine the effects of various components within Bella. Our analysis includes an exploration of the training costs associated with different ranks and their consequent influence on model performance. Given that Bella incorporates multiple parameter particles, we also delve into how varying the number of these particles affects Bella’s efficacy. Additionally, we explore the application of low-rank adapters across different layers and assess their impact. Further details on other studies are in Appendix E. We show in Table 7 that we achieve state-of-the-art performance on CAMELYON17.
Ablations on rank $r$As outlined in Section 4, we substitute the network’s extensive full matrix with low-rank matrices defined by the ‘rank’ parameter ($r$). This section aims to assess how this parameter influences Bella’s performance.
To thoroughly assess the influence of rank, we conduct our analysis on the challenging large-scale CAMELYON17 dataset. The results, depicted in Figure 5 (left), elucidate the relationship between rank size and performance. While utilizing a smaller rank considerably reduces the parameter space, it also restricts Bella’s learning capacity, hindering its ability to achieve optimal performance. In contrast, increasing the rank to 4 significantly enhances Bella’s efficiency. However, performance tends to plateau at a rank of 16, indicating a saturation point.
Figure 5: The impact of ranks on CAMELYON17 performance (left), as well as the impact of the number of parameter particles on CIFAR-10 (right) on Bella performance.
Ablations on the number of particles $n$In this section, we delve into the influence of the quantity of parameter particles on the performance of Bella. The findings, depicted in the right in Figure 5, reveal an improvement in Bella’s performance with an increase in the number of particles.
Datasets Metrics Bella Models Baseline Models (Base)
Ens. SVGD Single VI SGLD Ens. SVGD
CIFAR10 ECE $\downarrow$ 1.4 1 8.1 0.99 0.94 2.5 0.55
MCE $\downarrow$ 65 26 35 33 24 23 21
Brier $\downarrow$ 0.40 0.32 0.49 0.43 0.36 0.49 0.33
AUROC $\uparrow$ 99.98 99.99 99.78 99.93 99.94 99.99 99.99
CIFAR100 ECE $\downarrow$ 5.2 4.9 6.5 4.2 3.8 5.1 3.3
MCE $\downarrow$ 19 15 86 47 41 48 46
Brier $\downarrow$ 0.23 0.19 0.23 0.2 0.19 0.19 0.18
AUROC $\uparrow$ 99.88 99.88 99.71 99.87 99.83 99.88 99.88
CAMELYON17 ECE $\downarrow$ 5.5 5.3 6.4 5 5.2 5.3 4.9
MCE $\downarrow$ 31 29 34 36 31 30 21
Brier $\downarrow$ 5.2 5.1 5.5 5.4 5.4 5.1 4.5
AUROC $\uparrow$ 98.4 98.7 98.12 98.6 98.6 98.5 98.81
DomainNet ECE $\downarrow$ 5.9 5.8 7.9 5.1 4.6 5.2 4.9
MCE $\downarrow$ 40 37 61 40 41 37 34
Brier $\downarrow$ 5.3 5.2 6.3 5 4.4 5 4.7
AUROC $\uparrow$ 98.31 98.46 98.1 98.99 99.21 98.88 99.01
Table 3: Calibration comparison of Bella models (SVGD and Ensemble) with baseline methods across multiple datasets. The evaluation considers calibration metrics (ECE, MCE) and discriminative capability (Brier Score, AUROC). Lower values for ECE, MCE, and Brier Score, along with higher values for AUROC, indicate better performance.
The plot in Figure 2 (right) for the CAMELYON17 dataset begins with Bella at $r=2$ (comprising $5$ particles) and concludes with full SVGD (SVGD base) using the same number of particles. As $r$ increases, the accuracy of Bellas improves, eventually plateauing at approximately 95.08%, which is comparable to the accuracy of full SVGD (95.21%). This highlights the efficiency and advantages of our proposed Bella. Remarkably, Bella achieves performance on par with full SVGD models despite utilizing a significantly smaller pool of trainable parameters—approximately 0.3% for $r=4$—compared to the more parameter-intensive alternatives.
Calibration Study. Table 3 presents a comparison of the Bella models (SVGD and ensemble) against baseline models, including Variational Inference (VI) (Kim and Hospedales 2023) and Stochastic-Gradient Langevin Dynamic (SGLD) (Welling and Teh 2011), using several key metrics: Expected Calibration Error (ECE), Maximum Calibration Error (MCE), Brier score, and Area Under the Receiver Operating Characteristic curve (AUROC). Lower ECE and MCE values indicate that the model’s probability estimates are more reliable. A lower Brier score reflects better accuracy in the model’s probability predictions, while a higher AUROC demonstrates superior discrimination between classes. Across the datasets, the Bella models demonstrate performance comparable to the baseline models. This suggests that the Bella models, while utilizing far fewer parameters for training, offer a strong alternative to traditional baseline approaches while maintaining competitive calibration and Discriminative capability.
5.7 Generalization to a Visual Question Answer
In this section, we extend the application of our Bella to another challenging vision task, Visual Question Answering (VQA), as detailed in (Antol et al. 2015). We leverage the state-of-the-art, pre-trained, large multi-modal model LlaVA  (Liu et al. 2023b, a) for this purpose. Utilizing LlaVA transforms VQA into a process where an image and a natural-language question are inputs, and the model generates a free-form, open-ended text answer. Answering questions in VQA requires various intelligent capabilities, including not only image recognition but also complex reasoning. For this task, we employ VQA v2 (Antol et al. 2015) dataset containing 204,721 images, more than 1 Billion (1B) questions and 10B ground-truth answers in total. There are three main types of answers: Yes/No, a Number, and Other.
Model. In our experiments, we employed our proposed Bella on top of LlaVA-1.5-7B (Liu et al. 2023b, a). Further details about the dataset, model and metrics are deferred to Appendix H.
Accuracy.  The primary outcomes for Yes/No and Number answer queries are detailed in Table 4, respectively. We chose these question types to ensure a fair comparison and avoid semantic mismatches in open-ended answers in Others. A distinctive feature of Bella is its reduced uncertainty estimates for currect predictions, coupled with increased uncertainty estimates for incorrect ones. Moreover, it surpasses the Single base model regarding Accuracy and Exact Match metrics.
Notably, our method is efficient, particularly in contrast to the resource-intesive baseline Bayesian models tailored for this task (e.g. (Abbasnejad et al. 2019, 2020b, 2020a)). The computational requirements for using these baselines render the application of full SVGD or ensemble alternatives impractical, thereby highlighting the impact on practical applications by harnessing the effectiveness of Bella.
VQA: Evaluation of Yes/No Questions
Models Ent. Corrects ($\downarrow$) Ent. Incorrect ($\uparrow$) Accuracy ($\uparrow$) Match ($\uparrow$)
Single base 0.3336 0.5920 91.59 86.74
Ensemble Bella 0.3438 0.5935 91.20 86.21
SVGD Bella 0.3245 0.5950 92.46 87.83
VQA: Evaluation of Number Questions
Single base 0.4148 0.9911 58.69 49.26
Ensemble Bella 0.4248 0.9929 57.86 48.68
SVGD Bella 0.4059 0.9816 60.19 50.99
Model Uncertainty and Human Confidence. Further, we investigate the relationship between model uncertainty attained from Bella and human confidence, facilitated by multiple annotators in the VQA dataset. For this, we measure the correlation between the model disagreement by measuring the predictive entropy vs the human annotations of the answers.
(a) Yes/No questions.
(b) Number questions.
In Figure 6 we shows the correlation between the entropy of the model outputs and human confidence levels. By setting a specific threshold for the model’s entropy, we effectively bifurcate our predictions into two distinct categories: those with lower entropy fall into the ‘Low Entropy’ group, signalling reduced uncertainty within the model’s assessments, while predictions surpassing this entropy threshold are allocated to the ‘High Entropy’ segment, indicative of greater uncertainty. Intriguingly, our observations reveal a consistent negative correlation between the model’s entropy levels and human confidence across different types of queries, such as Yes/No and Number questions. This pattern suggests that the model’s entropy can serve as a reliable gauge of certainty, mirroring human judgment in its response to varying levels of uncertainty.
6 Conclusion
A Practical Approach to Bayesian Neural Networks
Overview of Materials in the Appendices
  1. 1.
    Comparing diversity measures between SVGD Bella and its SVGD base model (Appendix A)
  2. 2.
  3. 3.
    Uncertainty measures on the corrupted CIFAR-10-C dataset between SVGD Bella models and SVGD models (Appendix C)
  4. 4.
    Generalization Performance of Models on CIFAR-10-C Task (Appendix D)
  5. 5.
    Additional comparison studies. Including SoTA results on (Challenge 2024) (Appendix E)
  6. 6.
    Impact of $\gamma$ parameter values on robustness of models (Appendix F)
  7. 7.
  8. 8.
    Detailed information about VQA task (Appendix H)
  9. 9.
    Detailed information regarding datasets and hyper-parameters utilized in the paper (Appendix I)
Appendix A Diversity Measures
Appendix B Generalization of CIFAR-10 Models on OOD dataset (STL-10)
In this section, we evaluate the generalization of networks trained on CIFAR-10 on a similar datasets, STL-10, as in (Miller et al. 2021).